### v2

# Balancing Intensity and Focality in Directional DBS Under Uncertainty: A Simulation Study of Electrode Optimization via Metaheuristic L1L1 Method

Abstract
Background and Objective: As Deep Brain Stimulation (DBS) technology advances toward directional leads and optimization-based current steering, this study aims to improve the selection of electrode contact configurations using the recently developed metaheuristic L1-norm regularized L1-norm fitting (L1L1) method. The focus is in particular on L1L1’s capability to incorporate a priori lead field uncertainty, offering a potential advantage over conventional approaches that do not account for such variability.
Methods: Our optimization framework incorporates uncertainty by constraining the solution space based on lead field attenuation. This reflects physiological expectations about the volume of tissue activated (VTA) and serves to avoid overfitting. By applying this method to 8- and 40-contact electrode configurations, we optimize current distributions within a discretized finite element (FE) model, focusing on the lead field’s characteristics. The model accounts for uncertainty through these explicit constraints, enhancing the feasibility, focality, and robustness of the resulting solutions.
Results: The L1L1 method was validated through a series of numerical experiments using both noiseless and noisy lead fields, where the noise level was selected to reflect attenuation within VTA. It successfully fits and regularizes the current distribution across target structures, with hyperparameter optimization extracting either bipolar or multipolar electrode configurations. These configurations aim at maximizing focused current density or prioritize a high-gain field ratio in a discretized FE model. Compared to traditional methods, the L1L1 approach showed competitive performance in concentrating stimulation within the target region while minimizing unintended current spread, particularly under noisy conditions.
Conclusions: The L1L1 method provides a valuable tool for assisting specialists in optimizing electrode configurations for DBS. By incorporating uncertainty directly into the optimization process, we obtain a noise-robust framework for current steering, allowing for variations in lead field models and simulation parameters.
keywords:
Deep Brain Stimulation (DBS) , Anterior Nucleus of Thalamus , Convex Optimization , Metaheuristics
[…]
1 Introduction
Deep Brain Stimulation (DBS) [volkmann2002introduction, Perlmutter_2006_DBS, Gardner_2013_history_DBS] is a neuromodulatory therapy used to alleviate symptoms in patients with movement and neurological disorders, including, e.g., Parkinson’s disease [rodriguez2005bilateral, picillo2016programming, krack2019deep, steffen2020bipolar], essential tremor [benabid1996chronic], refractory epilepsy [LEHTIMAKI_2016_DBS_targetting, Jarvenpaa_2020_Improving_ANTDBS, fasano2021experience], and dystonia [johnson2013neuromodulation], by implanting an electrode lead into a targeted subcortical region and delivering controlled electrical stimulation through its contacts to modulate pathological neural activity. The central objective when determining this region of axonal activation is to shape the resulting volume of tissue activated (VTA), elicited by activating specific electrode contacts along the lead, so that it aligns with the intended therapeutic target [mcintyre_2004_cellular, butson_2006_electrodesign, butson_2007_patient, miocinovic_2006_subthalamic].
Computational representations of DBS activation patterns are often categorized by the number and polarity of active contacts. In monopolar configurations, one or more cathodic contacts are driven against the distant implantable pulse generator (IPG), which serves as a reference at infinity [volkmann_basic_algorithm_DBS_2006]. Bipolar configurations confine both anode and cathode to contacts on the lead, creating a locally circulating current and producing a more spatially restricted field [steffen2020bipolar, DELI_2011_UnipolarBipolarComparison]. Multipolar configurations further extend the programmable space by allowing multiple contacts with mixed polarities to shape the field in a more flexible manner. As contemporary leads incorporate segmented or high-density arrays, the resulting combinatorial growth in possible patterns renders exhaustive manual programming infeasible [Anderson_2018].
Effective steering of the VTA depends on a combination of tissue properties [butson_2005_tissue, McIntyre_2006_OptimizingDBS], electrode geometry and orientation [willsie2015computational], and the anatomical variability of the target region [wei_2005_currentdensityDBS, van2015novel]. Clinical investigations such as [contarino2014directional, pollo2014directional, dembek2017directional] and the synthesis in [Steigerwald_dDBS_2019] demonstrated directional steering empirically through stimulation mapping. Complementary computational studies, including [alonso2016investigation, zhang2020steering], examined how geometry, contact orientation, and heterogeneous tissue conductivities influence field distribution and VTA through finite element modeling. Although [Anderson_2018] illustrates how convex formulations can be incorporated into directional lead control, it represents one contribution within a broader set of emerging model-based strategies aimed at systematically optimizing stimulation patterns for directional DBS.
We employ the metaheuristic L1-norm regularized L1-norm fitting (L1L1) method [galazprieto_2022_L1vsL2] to retrieve electrode configurations over standard 8-contact and a complex 40-contacts directional DBS leads. The L1L1 method treats both the control vector and the data misfit with L1 penalties, which makes the optimization resistant to noise and model mismatch while discouraging diffuse or overly complex solutions, unlike least-squares formulations. We include configurations following Reciprocity Principle (RP2) [FERNANDEZCORAZZA2020116403], which leverages reciprocity theory to infer optimal stimulation patterns, and Tikhonov Regularized Least Squares (TLS) [dmochowski2011optimized], which balances data fidelity and solution smoothness through L2-norm regularization. For clarification: the terms “L1” and “L2” refer to the discrete 1- and 2-norms, respectively, of finite-dimensional vectors. The capitalized form L1L1 is used solely as the algorithm name and does not refer to the Lebesgue functional spaces, i.e., $L^{1}$ or $L^{2}$.
2 Materials and Methods
The quasi-static forward problem was modeled using the Complete Electrode Model (CEM) formulation [pursiainen2017forward] and solved with the Finite Element Method (FEM). Electrical field distributions were interpolated over the volumetric mesh following [pursiainen2016, miinalainen2019]. All modeling, meshing, and optimization procedures were performed using the Zeffiro Interface (ZI) version 6.1.0 [pursiainen2025zeffiro], an open-source MATLAB-based pipeline that provides unstructured boundary-fitted, multi-compartment tetrahedral FE mesh generation from MRI-derived segmentations.
2.1 Realistic Head Model
We generated a tetrahedral-based mesh with a nominal element size of 3.0 mm from an open access T1 weighted MRI data set from a neurologically healthy adult [piastra_maria_carla_2020_3888381]. We performed cortical segmentations using FreeSurfer [fischl2012freesurfer], supported by functions from SPM12 package [ASHBURNER20091163]. Deep nuclei labels were extracted according to the Aseg atlas (Fig. 1(a)) resulting in a realistic head model with 18 anatomically distinct tissue compartments. A double-layered cylindrical compartment with a 1.27 mm outer diameter, and a 0.5 mm thick encapsulation layer [anderson_2019_anodic] was further added and interpolated using in-house scripts. The cylinder functions as a proxy for setting the electrodes with their respective geometric properties.
As our target sub-compartment would be then the anterior nucleus of the thalamus (ANT), consistent with ANT-DBS applications in epilepsy [LEHTIMAKI_2016_DBS_targetting, Jarvenpaa_2020_Improving_ANTDBS], the thalamus compartment underwent three levels of nested refinements subdividing every tetrahedron on each step into four congruent sub-elements [galazprieto_2023_mesh]. This procedure yielded element diameters ranging from 0.1725 to 0.375 mm in the refined region and produced a final mesh comprising approximately 4.4 million nodes and 25.4 million tetrahedral elements (Fig. 1(b)).
We assigned electrical conductivities as follows: 0.14 S/m for white matter, 0.33 S/m for gray matter, 0.0064 S/m for skull, 1.79 S/m for cerebrospinal fluid (CSF), 0.33 S/m for subcortical structures, and 0.33 S/m for skin, as reported in [dannhauer2011modeling]. Subcortical nuclei were assumed to share electrical conductivity values of gray matter, while ventricular regions were assigned the conductivity of CSF, in accordance with [rezaei2021reconstructing] and its referenced sources.
2.2 Lead Field Matrix
To establish a baseline representation for spatial attenuation of the lead field and its correspondence to feasible VTA coverage, we derived an initial low-resolution lead field (LR-LF) comprising $244\times 3$ dipolar DOFs distributed throughout the thalamus. The spatial extent for detailed current steering was then guided by reported activation radii around the electrodes, where typical stimulation produces tissue activation on the order of approximately 2.0–4.0 mm [butson_2007_patient, miocinovic_2006_subthalamic], and may extend to roughly 3.8–4.0 mm under clinically relevant parameters [madler2012activated]. We then establish a spherical region of 6.0 mm radius centered on the lead to represent the domain in which fine-grained modulation is physiologically meaningful. A high-resolution lead field (HR-LF) was subsequently generated by increasing the dipole density to $3620\times 3$ DOFs across the same anatomical region and then retaining only those sources falling within the 6.0 mm domain (Fig. 1(c)). A full derivation can be found in A.1.
(a) Segmented head model domain depicting main subcortical structures.
(b) Refined unstructured boundary-fitted thalamus sub-compartment with activity (smoothed).
(c) Superior, coronal, and sagittal view of target domain. The interthalamic adhesion can be easily noticed.
(d) 8- and 40-electrode contacts, respectively.
Figure 1: (1(a)): A head model including skull, forebrain, midbrain, hindbrain structures as well as a directional deep brain stimulation (DBS) lead positioned at the right-hemispheric anterior nucleus of thalamus (ANT). (1(b)): Refined unstructured boundary-fitted thalamus sub-compartment with the distribution of tetrahedral mesh edges and an interpolated current intensity distribution shown on the surface. (1(c)): Placement of a 40-contact dDBS lead with a bipolar electrode configuration (anodal as red, and cathodal as blue). The dipolar target current (red) is depicted by the red pin, indicating the target position and orientation; a 6.0 millimeter ($m$) fitting region is displayed as an arrangement of source points (yellow squares) for High-Resolution Lead Field (HR-LF) matrix. (1(d)): Probing strategy of the 8- and 40-electrode contacts inspired by Abbott Infinity™ and Medtronic-Sapiens, respectively.
2.3 Optimization Framework
The applied optimization framework aims to return a focal anodal–cathodal electrode configuration that best reproduces the volumetric current distribution associated with a prescribed target current dipole model [demunck1988, mosher2002multiple, medani_2012_ModellingDipoles]. The framework enhances field components aligned with the target while attenuating directions dominated by nuisance activation. Active electrodes are selected iteratively by adjusting the current pattern $\mathbf{y}$ under a set of predefined hyperparameters and an expected number of active contacts $K$, depending on the chosen algorithm (e.g., Reciprocity Principle, L1L1 Method, or Tikhonov-regularized least squares). Metaheuristics are employed to explore the feasible design space and identify current patterns that best satisfy the required performance characteristics [galazprieto_2024_lattice].
To restrict the number of admissible solutions, we impose a lower bound on the focused current density $\Gamma$ ($\Gamma\geq\Gamma_{0}=0.80~\mathrm{mA}$) to ensure sufficient activation at the target, as well as an upper bound on the nuisance current density $\Xi$ to limit excessive off-target stimulation. The candidate solution is thus defined as the current pattern that maximizes the field ratio between target- and nuisance-region current densities while satisfying these constraints.
As safety constraints, the total (absolute) injected current is constrained to $\mu=4.0~\mathrm{mA}$ (6), with per-contact limits $\gamma=\pm 2.0~\mathrm{mA}$ (7). Electrode potentials are normalized to satisfy the net-current constraint (8). Lastly, the stimulation is, however, treated as time-invariant, as the objective of this work is to characterize the spatial organization of current flow and the resulting current density distributions, rather than to model pulse dynamics or frequency-dependent neuronal responses.
For reference magnitude $\|{\bf x}_{1}\|_{2}=3.85$ A/m2 corresponds to the activation threshold reference. This value is obtained by dividing the average dipole moment density of pyramidal neurons (0.77 nAm/mm2) by an approximate dendritic axis length of 0.2 mm [MURAKAMI201549].
2.3.1 Electrode Configurations
[…]
2.3.2 Bipolar Configurations
Bipolar patterns are obtained using the reciprocity principle (RP) [FERNANDEZCORAZZA2020116403]. The principle states that the maximal focused current density $\Gamma_{\mathrm{max}}$ for a linear quasi-static system is achieved by selecting a single anode–cathode pair. In practice, the target dipolar field is multiplied by the transpose of the lead field matrix, and the contacts corresponding to the strongest positive and negative entries are chosen as anode and cathode, respectively, with all remaining contacts set to zero (Fig. 2). RP maximizes the delivered current density at the target but does not optimize the ratio between target activation and nuisance activation, and thus may yield suboptimal field selectivity.
2.3.3 Multipolar Configurations
Multipolar configurations are derived using the L1-norm regularized L1-norm fitting (L1L1) method [galazprieto_2022_L1vsL2] and Tikhonov-regularized least squares (TLS) [dmochowski2011optimized]. Both methods assume a predetermined set of candidate active contacts and estimate the corresponding stimulation amplitudes under polarity, charge, and safety constraints (detailed in A.2).
The L1L1 formulation imposes an L1 penalty on both the stimulation vector and the residual, which enhances robustness to lead-field uncertainty and promotes sparse, interpretable support. TLS, in contrast, employs an L2 penalty that stabilizes the solution in the presence of correlated columns and numerical ill-conditioning. Because the full combinatorial search space of possible contact subsets grows exponentially with $K$, both formulations employ a metaheuristic support-selection strategy to efficiently explore only a tractable subset of candidate configurations while maintaining practical computational cost [galazprieto_2024_lattice].
8-contact Electrode Configuration
Refer to caption RPRefer to captionRefer to caption Refer to caption L1L1(A)Refer to captionRefer to caption Refer to caption L1L1(B)Refer to captionRefer to caption Refer to caption TLSRefer to captionRefer to caption
40-contact Electrode Configuration
Refer to caption RPRefer to captionRefer to caption Refer to caption L1L1(A)Refer to captionRefer to caption Refer to caption L1L1(B)Refer to captionRefer to caption Refer to caption TLSRefer to captionRefer to caption
Figure 2: Examples of current injection patterns obtained using the Reciprocity Principle (RP), metaheuristic L1-norm regularized L1-norm fitting (L1L1) method under two different nuisance current thresholds (L1L1(A) with $\varepsilon=\{0:-160\}$ and L1L1(B) with $\varepsilon=\{0:-10\}$), and the Tikhonov-Regularized Least Square (TLS) method. The target position and orientation (parallel to the lead) is indicated by the red upward pin. The bar plots display the resulting electrode current amplitudes, while the contact maps visualize their spatial distribution over the electrodes; cathodal (blue), anodal (red), and inactive (gray/green). The RP method yields a bipolar configuration maximizing focused current density $\Gamma$, while the remaining methods searches for a multipolar configuration maximizing focality $\Theta=\Gamma/\Xi$ under an a priori constraint $\varepsilon$ on the nuisance current density $\Xi$, determined by the peak signal-to-noise ratio (PSNR) $\delta^{-1}$ and the lead field attenuation. The L1L1(B) result reflects more strongly constrained $\Xi$ (smaller dynamic range), activating fewer contacts whereas L1L1(A) and TLS allow higher $\Theta$ values and broader excitation patterns. The voltage/current on the electrode contacts are $\pm 2.0$ mA (milliamperes).
2.4 Uncertainty Modeling
Uncertainty refers to various physiological and technical factors that affect the accuracy and reliability of the stimulation, including imprecision in lead placement, patient-specific anatomical variation, and inhomogeneities in tissue conductivity [schmidt2016uncertainty, athawale2019statistical]. Potential sources affecting the optimization framework are modeled as additive perturbations to the lead field matrix through Gaussian noise model. While this is a simplification, it enables a tractable first-step analysis of robustness. We acknowledge that in practice, the true uncertainty arises from a variety of unrelated sources.
We postulate that the volumetric current density distribution can be optimized for DOFs indexed by $i$ whose corresponding reduced lead field row ${\bf\tilde{l}}_{i}$ has an L2-norm above a threshold. This row is first reduced and zero-averaged, taking the form
$\tilde{\bf l}_{i}=\left(L_{i,1}-\overline{L}_{i},L_{i,2}-\overline{L}_{i},\dots,L_{i,K}-\overline{L}_{i}\right),$
where $\overline{L}_{i}=K^{-1}\sum_{j=1}^{K}L_{i,j}$ is the mean of the row entries. The threshold is determined via a peak signal-to-noise ratio (PSNR) constraint, ensuring that only DOFs contributing significantly to the lead field are included in the optimization. The feasible set of DOFs is defined as
$\mathcal{V}_{\delta}=\{i\,|\,\|{\bf\tilde{l}}_{i}\|_{2}\geq\delta\max_{k}\|{\bf\tilde{l}}_{k}\|_{2}\},$
where $\max_{k}\|{\bf\tilde{l}}_{k}\|_{2}$ is the largest L2-norm among all the zero-averaged rows in the lead field rows. The parameter $\delta$ sets a relative threshold, filtering out DOFs whose contributions are too weak to be meaningful. When $\delta$ is interpreted as the activation threshold, $\mathcal{V}_{\delta}$ approximates the maximum VTA (Fig. 3(a)).
Because lead field attenuation and uncertainty impose limits on stimulation accuracy, the PSNR constraint effectively defines the theoretical performance limit of the optimization. If the target site corresponds to index $\ell$, the maximum achievable dynamic range $\varepsilon^{-1}$ between the current density at the target and surrounding regions is approximately
$\varepsilon\geq\delta\,\|{\bf\tilde{l}}_{\ell}\|_{2}^{-1}\max_{k}\|{\bf\tilde{l}}_{k}\|_{2}.$
This indicates that the minimum contrast between the target and non-target regions is constrained by both $\delta$ and the structure of the lead field matrix (Fig. 3(b)).
[…]
(a) Left: The set $\mathcal{V}_{\delta}$ corresponding to PSNR $\delta^{-1}$ following from the uncertainty factors is shown by the blue dashed circle with radius $d_{\delta}$. The target is at the distance $d_{T}$ from the probe. Between the target and the boundary of $\mathcal{V}_{\delta}$ is a range $d_{\varepsilon}$ in which the lead field decays by the factor $\varepsilon$, which approximates the maximum achievable dynamic range $\varepsilon$ between the focused current density at the target position and in its surroundings. Right: One-dimensional interpretation of the lead field decay.
$\mathcal{V}_{-10\text{ dB}}$
$\mathcal{V}_{-20\text{ dB}}$
$\mathcal{V}_{-30\text{ dB}}$
$\mathcal{V}_{-40\text{ dB}}$
(b) The set $\mathcal{V}_{\delta}$ illustrated for the volumetric model of this study with respect to PSNR $\delta^{-1}$ levels from 10 to 40 dB.
Figure 3: (3(a)): A schematic of how the peak signal-to-noise ratio (PSNR) $\delta^{-1}$ relates to the set $\mathcal{V}_{\delta}$ and to the approximate maximum dynamic range $\varepsilon^{-1}$ of the volumetric stimulation current field. (3(b)): The set $\mathcal{V}_{\delta}$ shown for PSNR 10–40 dB, providing a rough approximation of the maximum feasible coverage of volume of tissue activated (VTA) for an activation threshold set by $\delta$.
2.5 Hardware
Dell Precision 5820 Workstation equipped with 256 GB of RAM, a 10-core Intel i9–10900X CPU, and an NVidia Quadro RTX 4000 GPU. The GPU were used to accelerate both forward and inverse calculations.
2.6 Numerical Experiments
Our numerical experiments evaluate the proposed optimization methods with respect to: (i) 8- and 40-contact DBS lead configurations, (ii) noiseless and noisy forward mappings, (iii) dipole orientation relative to the lead array (parallel or perpendicular) [Janson_2020_Activation], and (iv) algorithm efficiency.
Table 1: Summary of the proposed optimization methods. RP provides a bipolar reciprocal of the target, whereas L1L1 and TLS perform an extensive metaheuristic search. L1L1 method is subdivided into L1L1(A) and L1L1(B) according to two alternative a priori assumptions for the feasible $\varepsilon$ threshold range: [-160, 0] dB and [-10, 0] dB, respectively. The regularization and L2-norm weighting parameter ranges are adopted from [galazprieto_2022_L1vsL2]. CPU time refers to the computation time needed to evaluate a single optimizer with fixed hyperparameters for the HR-LF using the 40-contact lead.
Hyperparameter CPU
Tuning Range Role Time
Method Search Type (dB) (s)
RP Reciprocal 5.7E-03
L1L1(A) Metaheuristic $\alpha\in[-100,-30]$ Regularization 6.0
$\varepsilon\in[-160,0]$ Threshold
L1L1(B) Metaheuristic $\alpha\in[-100,-30]$ Regularization 3.7
$\varepsilon\in[-10,0]$ Threshold
TLS Metaheuristic $\gamma\in[-200,-110]$ Regularization 5.0E-02
$\beta\in[-50,40]$ Weight
The L1L1 method is tested under two alternative a priori assumptions for the feasible $\varepsilon$ threshold range: [-160, 0] dB (L1L1(A)) and [-10, 0] dB (L1L1(B)). The first neglects uncertainty in the field ratio $\Theta$, while the second accounts for variability within a 10 dB dynamic range. TLS uses regularization parameter $\gamma$ and L2-norm weighting parameter $\beta$ intervals as in [galazprieto_2022_L1vsL2] (Table 1). Detailed formulation of both the L1L1 and TLS methods is provided in A.2 and A.3, respectively.
To examine sensitivity to measurement and modeling errors, Gaussian noise is added to the lead field. Candidate solutions are computed for each target DOF of the LR-LF under PSNR 40.0 dB corruption. While anatomical variations can introduce spatially correlated errors, this additive noise provides a tractable baseline to assess robustness to model perturbations. Noise sensitivity is specifically analyzed for two target positions (I) and (II) (Fig. 4) along the HR-LF using PSNRs of 40.0 and 50.0 dB. Interpreted in terms of $\mathcal{V}_{\delta}$, PSNR 40.0 dB (50.0 dB) corresponds to the assumption that each target in $\mathcal{V}_{-30\text{ dB}}$ ($\mathcal{V}_{-40\text{ dB}}$) can be stimulated with at least a 10 dB dynamic range $\varepsilon^{-1}$ for $\Theta$, which aligns with the $\varepsilon$ range of [-10, 0] used for L1L1(B).
Figure 4: Visualization of the two synthetic stimulation targets, Target (I) and Target (II), analyzed separately in numerical experiments. Both targets are positioned within the right thalamus near a directional DBS lead. The figure also displays the lead’s center point and a concentric spherical region of interest (ROI) with 6 mm radius, representing an approximate VTA. This ROI defines the region where maximizing stimulation intensity while minimizing current spread to surrounding areas is considered feasible. As targets (I) and (II) are not right next to the probe, lead field uncertainty affects the optimization process. Bilateral thalamic structures are shown in gray for anatomical reference. The distances are shown with respect to the boundary cylindrical probe illustrated by the ring around the center point.
LR-LF HR-LF
Focused $\Gamma$ (A/m2) Nuisance $\Xi$ (A/m2) Ratio $\Theta$ (rel.) Focused $\Gamma$ (A/m2) Nuisance $\Xi$ (A/m2) Ratio $\Theta$ (rel.)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
Figure 5: Volumetric optimization results for the 8-contact probe and noiseless lead fields. Decision variable $\Gamma$, $\Xi$ and $\Theta$ for optimizers found via Reciprocity Principle (RP), L1-norm regularized L1-norm fitting (L1L1) evaluated as L1L1(A) and L1L1(B) with feasible $\varepsilon$ threshold range: [-160, 0] dB and [-10, 0] dB, respectively, and Tikhonov Regularized Least Squares (TLS) methods illustrated as a function of the target position and orientation (parallel or perpendicular w.r.t. the probe). Left: distributions for a Low-Resolution Lead Field (LR-LF) where the spatial degrees of freedom (DOFs) cover both thalamus lobes. Right: distribution for a High-Resolution Lead Field (HR-LF) where the DOFs is confined to a region of interest within a radius of 6.0 mm (millimeters) from the center of the lead. The optimal electrode configuration found by the L1L1 method maximizes the field ratio $\Theta$ subject to $\Gamma>=\Gamma_{0}$. The $\Theta$ distribution found using L1L1(A) is elevated as compared to the case of RP and TLS. The limited dynamic range ($\varepsilon$ range) of L1L1(B) is reflected in an overall slightly lower $\Theta$.
LR-LF HR-LF
Focused $\Gamma$ (A/m2) Nuisance $\Xi$ (A/m2) Ratio $\Theta$ (rel.) Focused $\Gamma$ (A/m2) Nuisance $\Xi$ (A/m2) Ratio $\Theta$ (rel.)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
Figure 6: Volumetric optimization results for the 40-contact probe and noiseless lead fields. As compared to the results obtained with 8-contact probe the overall level of $\Theta$ is greater for L1L1(A), TLS and RP, while for L1L1(B) with limited dynamic range ($\varepsilon$ range) there is no clear increase.
3 Results
The applied algorithms generate distinct distributions of the variables $\Gamma$, $\Xi$, and $\Theta$, which can be differentiated based on how they are spatially organized across the target domain under the optimal electrode configurations for the 8- (Fig. 5) and 40-contact (Fig. 6) leads. The direction of the current dipole plays the role in shaping electric field targeting. The anode–cathode pair exhibiting the strongest positive and negative values may activate electrodes located as close as $0.189\pm 0.009\,\text{mm}$ or as far as $6.375\pm 0.375\,\text{mm}$ apart across the lead. With the parallel orientation, the targeting is highly concentrated near the tip of the probe, whereas the perpendicular orientation results in a broader spatial spread, with activity shifted medially relative to the lead.
Bipolar configurations obtained via RP yield the highest focused current density $\Gamma$; however, this comes at the expense of focality, as indicated by elevated nuisance current $\Xi$ and a reduced field ratio $\Theta$. This trade-off is understandable, as the multipolar configurations are designed to maximize the field ratio $\Theta$ subject to the constraint $\Gamma\geq\Gamma_{0}>0$, with $\Gamma_{0}=0.8\,\text{mA}$, thereby allowing activation of more than two contacts.
Figure 7 and Table 2 compare noiseless and noisy ($\text{PSNR}=40\,\text{dB}$) results for $\Gamma$ and $\Theta$ using the L1L1 method over a LR-LF. In the noiseless setting, the L1L1(B) reconstructions with $\varepsilon\in[-10,0]\,\text{dB}$ yield slightly higher $\Gamma$ and $\Xi$ values, and moderate $\Theta$ compared to L1L1(A) with $\varepsilon\in[-160,0]\,\text{dB}$. L1L1(A) provided unrealistically high values of the field ratio $\Theta$: up to $27.07$ in the 8-contact directional DBS leads and as high as $536.05$ in the 40-contact case indicating a case of overfitting.
Figure 8 further examines the noise sensitivity of the applied optimization methods using two specific target positions, illustrated in Figure 4, covering a sample of twenty different noise realizations for both PSNR 40 and 50 dB. The results demonstrate that in a noisy situation, L1L1(A) and L1L1(B) provide similar distributions for $\Gamma$ and $\Theta$. Compared to RP and TLS, the L1L1 approach seems to provide elevated field ratio $\Theta$ with a difference that is only marginal or absent with PSNR 40 dB, but observable with PSNR 50 dB. Accordingly, the potential benefit of using L1L1 depends critically on the level of modeling uncertainty, and it vanishes when this uncertainty becomes excessive. Of the four optimization approaches tested, L1L1(B) provides the closest agreement between noiseless and noisy outcomes and proves particularly effective in maximizing the field ratio $\Theta$ in the presence of noise.
Table 2: Maximum focused current density and field ratio described as $\Gamma_{max}$ and $\Theta_{max}$, respectively, for a 8- and 40-contacts multipolar DBS lead using the L1L1 method over a low-resolution lead field (LR-LF), with current dipole targeting aligned parallel and perpendicular to the DBS lead, for both noiseless (No Noise) and noisy (PSNR 40 dB) settings.
8-contacts Multipolar
L1L1(A), $\varepsilon\in[-160,0]$ L1L1(B), $\varepsilon\in[-10,0]$
DV Orientation No Noise PSNR 40 dB No Noise PSNR 40 dB
$\Gamma_{max}$ Parallel 2.42 2.27 2.27 2.27
Perpendicular 3.04 2.35 2.28 2.34
$\Theta_{max}$ Parallel 21.53 5.56 13.08 5.48
Perpendicular 27.07 7.49 13.18 7.22
40-contacts Multipolar
L1L1(A), $\varepsilon\in[-160,0]$ L1L1(B), $\varepsilon\in[-10,0]$
DV Orientation No Noise PSNR 40 dB No Noise PSNR 40 dB
$\Gamma_{max}$ Parallel 2.20 1.74 2.02 2.24
Perpendicular 2.24 2.29 3.37 2.23
$\Theta_{max}$ Parallel 526.44 7.56 9.17 8.49
Perpendicular 536.05 9.96 10.97 8.47
8-contact Electrode Configuration
Noiseless lead field
A1. Focus $\Gamma$ (A/m2)
B1. Ratio $\Theta$ (rel.)
Noisy lead field
C1. Focus $\Gamma$ (A/m2)
D1. Ratio $\Theta$ (rel.)
40-contact Electrode Configuration
Noiseless lead field
A2. Focus $\Gamma$ (A/m2)
B2. Ratio $\Theta$ (rel.)
Noisy lead field
C2. Focus $\Gamma$ (A/m2)
D2. Ratio $\Theta$ (rel.)
L1L1(A), $\varepsilon\in[-160,0]$, Parallel
L1L1(B), $\varepsilon\in[-10,0]$, Parallel
L1L1(A), $\varepsilon\in[-160,0]$, Perpendicular
L1L1(B), $\varepsilon\in[-10,0]$, Perpendicular
Figure 7: Performance of L1L1(A) and L1L1(B) for optimizing the field ratio $\Theta$ subject to criterion $\Gamma\geq\Gamma_{0}$ with noiseless and noisy (PSNR 40 dB) lead field. L1L1(A) and L1L1(B) suppress the nuisance current density $\Xi$ pattern using two different dynamic ranges, namely, $\varepsilon\in[-160,0]$ (blue) and $\varepsilon\in[-10,0]$ (red), respectively. The results are sub-grouped by the direction of the current dipole, whether its alignment is parallel (solid line) or perpendicular (dotted line) with the applied lead. The degrees of freedom have been sorted according to ascending field ratio $\Theta$. The results show that the overly optimistic dynamic range applied in L1L1(A) results in very high field values of $\Theta$ in the noiseless case, which are, however, absent in the noisy one. In contrast, L1L1(B) is only marginally affected by the noise due to more realistic assumption of the maximum obtainable dynamic range under uncertainty.
8-contact Electrode Configuration
Parallel
Focus $\Gamma$ (A/m2)
Ratio $\Theta$ (rel.)
Perpendicular
Focus $\Gamma$ (A/m2)
Ratio $\Theta$ (rel.)
Target (I)
Parallel
Focus $\Gamma$ (A/m2)
Ratio $\Theta$ (rel.)
Perpendicular
Focus $\Gamma$ (A/m2)
Ratio $\Theta$ (rel.)
Target (II)
40-contact Electrode Configuration
Parallel
Focus $\Gamma$ (A/m2)
Ratio $\Theta$ (rel.)
Perpendicular
Focus $\Gamma$ (A/m2)
Ratio $\Theta$ (rel.)
Target (I)
Parallel
Focus $\Gamma$ (A/m2)
Ratio $\Theta$ (rel.)
Perpendicular
Focus $\Gamma$ (A/m2)
Ratio $\Theta$ (rel.)
Target (II)
Figure 8: Boxplots showing the focused current density $\Gamma$ and field ratio $\Theta$ obtained with a noise-corrupted HR-LF and the 8-contact electrode configuration for test targets (I) and (II). The optimizers have been found through RP, L1L1(A), L1L1(B), TLS methods. Each bar collates a sample of twenty decision variable values corresponding to optimizers obtained with as many independent lead field noise realizations (PSNR 40 or 50 dB). The box shows the interquartile range (IQR) between the first quantile $Q1$ and the third one $Q3$, with a horizontal line at the median. The whiskers extend from the edges of the box to the smallest and largest values within $Q1-1.5\times\text{IQR},Q3+1.5\times\text{IQR}]$, where $\text{IQR}=Q3-Q1$. Data points outside this range are considered outliers and are plotted individually as markers. The results obtained with L1L1(A) and L1L1(B) demonstrate a comparable sensitivity under lead field noise, suggesting that the limited dynamic range of L1L1(B) does not downgrade the L1L1 method’s performance in situations involving uncertainty. Furthermore, both L1L1(A) and L1L1(B) find an overall elevated field ratio compared to RP and TLS with the greater PSNR level 50 dB, which allows for an improved +10 dB dynamic range for the optimization algorithm as compared to PSNR 40 dB.
4 Discussion
This study explores the application of the recently introduced metaheuristic L1L1 method [galazprieto_2022_L1vsL2] to optimize current patterns and electrode configurations for DBS leads. The L1L1 method was applied in two variants, L1L1(A) constrains $\varepsilon$ to the interval $160,0]\,\text{dB}$ and L1L1(B) to $10,0]\,\text{dB}$ reflecting different levels of a priori modeling uncertainty [schmidt2016uncertainty, athawale2019statistical]. For comparison, optimization performance using Reciprocity Principle (RP)  [FERNANDEZCORAZZA2020116403] and Tikhonov Regularized Least Squares (TLS) [dmochowski2011optimized] methods were included. The core advantage in applying the L1L1 method is the capacity to incorporate a priori uncertainty through an explicit constraint, offering an advantage over traditional approaches, such as the least-square methods, which often do not account for such variability.
Within the optimization framework, the focused current density $\Gamma$, nuisance current density $\Xi$, and field ratio $\Theta=\Gamma/\Xi$ are implemented as the three metrics to measure spatial current distributions influenced by the quality of the forward modelling, position and orientation of the current dipole, and uncertainness which is defined through a threshold on the nuisance current density, determined by the relative attenuation of the lead field matrix with respect to a target stimulation site. This attenuation is quantified in decibels and corresponds to typical falloff patterns observed in the volume of tissue activated (VTA) [zhang2020steering].
In noiseless conditions, solutions using a more permissive threshold value (L1L1(A) case) often exhibited unrealistically high field ratio values, approximately fifty times greater than those derived using a more conservative threshold (L1L1(B) case). The addition of noise promoted the optimizer’s reliance from convergence towards over-idealized solutions, forcing the algorithm to select current patterns that remain effective even under slight model variations, producing more physiologically realistic distributions.
The estimates for the set $\mathcal{V}_{\delta}$, where the lead field amplitude attenuates by a factor $\delta$ for $\delta\in[-40,-10]\,\text{dB}$, align closely with established knowledge regarding VTA. For $\delta\approx-40\,\text{dB}$, the model predicts attenuation within a spatial region consistent with a VTA radius of approximately 4.0 mm, in line with Butson and McIntyre [butson_2007_patient], Miocinovic et al. [miocinovic_2006_subthalamic], and Mädler and Coenen [madler2012activated]. This suggests that the noise levels used (PSNR 40 dB and 50 dB) represent reasonable bounds for physiological uncertainty. At PSNR 40 dB, the dynamic range of $\Theta$ is determined by the maximum lead field attenuation within the VTA, whereas PSNR 50 dB supports the assumption that each position within the VTA can be independently and coherently targeted.
When comparing electrode configurations, increasing the number of contact points generally improves adaptability and control over the resulting electric field. Parallel configurations, i.e., alignment of the current dipole with the DBS lead, produced localized activation near the tip of the lead, whereas perpendicular configurations yielded a broader medially shifted targeting. This is driven by reduced distances between electrodes and target locations, better alignment of the induced field with the target current direction, and greater flexibility due to more DOFs.
Overall, the findings provide a proof-of-concept for integrating a priori uncertainty modeling into numerically simulated DBS optimization. When sources of uncertainty, such as anatomical variability, conductivity inhomogeneities, or lead placement deviations, are known or estimated, the L1L1 method can improve focality and directionality of stimulation, enabling more selective targeting of desired structures while minimizing off-target activation. Such improvements have the potential to enhance therapeutic precision and reduce side effects in clinical practice [fasano2021experience].
4.1 Limitations and Future Work
While the current uncertainty modeling strategy provides a principled framework for constraining the optimization based on lead field attenuation, it also has several limitations. First, the threshold values used to define acceptable nuisance current levels are heuristic and may not fully capture the complexity of individual patient anatomies or tissue conductivities. This simplification assumes spatial uniformity in attenuation, which may not hold, for example, in the presence of anisotropic or heterogeneous tissue properties [dannhauer2011modeling]. Additionally, the approach relies on a static surrogate model of the VTA and does not directly incorporate a probabilistic model of uncertainty [schmidt2016uncertainty, athawale2019statistical], which would allow for a more comprehensive treatment of variability. Potential alternative surrogate approaches exist, such as, threshold-based curve fitting [butson_2006_electrodesign], and statistical emulators such as Gaussian process classifiers [DeLaPava_2017_accelerating, Orozco_2019_KernelDBS], which approach VTA in different ways.
The Gaussian white-noise model is adopted because it furnishes the maximally uninformative representation of uncertainty under a finite-variance constraint [Jaynes2003, CoverThomas2006]. Given that the present optimization framework does not specify a detailed probabilistic law for perturbations, this choice avoids introducing unsupported structure and provides a neutral, information-theoretic baseline. More sophisticated uncertainty quantification techniques, such as Markov Chain Monte Carlo (MCMC)-based Bayesian sampling [kaipio2006statistical] or deep-learning approaches to predictive uncertainty [schmidt2016uncertainty, athawale2019statistical], could be employed when richer data are available.
Compared to the reference methods, the L1L1 method is also slightly more demanding in terms of computing time. However, the increase is not significant and does not limit its practical use. Other mathematical optimization techniques, such as TLS, may offer promising alternatives alongside the L1L1 method for further refining the a priori uncertainty modeling adopted in this study. A logical direction for future research is the development of a posteriori models, where the current approach could be validated in more advanced experimental or clinical settings. This would enable a deeper understanding of the potential hyperparameter realizations in real-world applications. Translating these computational findings into clinical practice will require rigorous validation through empirical testing and clinical trials.
This manuscript does not address clinical implications; however, for the configuration studied, certain anatomical structures, such as the lamina medullaris near the modeled lead trajectory, may require special consideration in clinical practice to avoid stimulation-induced side effects. Translating these methodological results to clinical applications will necessitate examining the relationship between the L1L1 method and uncertainties in individual factors such as lead placement, tissue conductivity and anisotropy, and patient-specific anatomical variations including thalamic segmentation. To support this, we employ finite element (FE) discretizations incorporating thalamic subdivisions through a boundary-fitted multi-compartment FE mesh together with a Complete Electrode Model (CEM) [pursiainen2017forward], as in the present study.
5 Conclusions
This study represents an advancement in optimizing the volumetric current fields for DBS by leveraging the metaheuristic L1L1 method. Previous contact selection studies, such as [Anderson_2018] and later [Janson_2020_Activation], have used convex optimization to determine optimal stimulation settings and this represents at addressing varying levels of a priori lead field uncertainty, which can be attributed as patient anatomy, misleading lead placement, and tissue properties, which often vary in clinical practice.
Our findings demonstrate that the method can achieve relatively high field ratios with a carefully chosen activation threshold $\varepsilon$ value. This, in turn, has the potential to improve individualized stimulation outcomes. An ongoing work is to integrate patient-specific data, such as thalamic segmentation, to improve stimulation accuracy. This involves using boundary-fitted multi-compartment FE meshes [galazprieto_2023_mesh] and a Complete Electrode Model [pursiainen2017forward], as in this study.
6 Conflict of Interest
The authors certify that this study is a result of purely academic, open, and independent research. They have no affiliations with or involvement in any organization or entity with a financial interest or non-financial interest, such as personal or professional relationships, affiliations, knowledge, or beliefs in the subject matter or materials discussed in this manuscript.
7 Author Contributions
FGP: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Visualization, Writing—original draft. AL: Conceptualization, Data curation, Investigation, Methodology, Project administration, Writing—original draft. MS: Formal analysis, Supervision, Validation, Writing—review & editing. SP: Conceptualization, Formal analysis, Funding acquisition, Investigation, Methodology, Project administration, Resources, Software, Supervision, Validation, Visualization, Writing—review & editing.
8 Funding
FGP, MS, and SP were supported by the Research Council of Finland (RCF) through the Center of Excellence in Inverse Modelling and Imaging 2018–2025 (RCF 353089), Exploratory Study for Radar Tomography of Dimorphos (RCF 359198), and the Flagship of Advanced Mathematics for Sensing, Imaging and Modelling (FAME) (RCF 359185); DAAD (German Academic Exchange Service) project ”Non-invasively reconstructing and inhibiting activity in focal epilepsy” (354976, 367453); and the ERA Personalised diagnosis and treatment for refractory focal paediatric and adult epilepsy (PerEpi) project (RCF 344712). MS holds Academy Research Fellowship 2025–2029 (RCF 371055).
9 Ethical Statement
This work is based exclusively on mathematical modeling and numerical simulations, using an openly available MRI dataset [piastra_maria_carla_2020_3888381] as input. No new data was collected, and no sensitive or personally identifiable information was processed at any stage of the study.
10 Acknowledgments
The authors would like to thank Prof. Dr. rer. nat. Carsten H. Wolters, researchers and clinicians from the Institute for Biomagnetism and Biosignalanalysis (IBB), and the PerEpi consortium for their continuous support, discussions, and feedback regarding non-invasive brain stimulation topics, regression analysis, and the seminars prepared during the elaboration of this study.
Appendix A Mathematical Optimization Scheme
A.1 Formulation of the Lead Field Matrix
The linear forward mapping between electrode currents and the resulting volumetric current density distribution is given by the real-valued $(N+M)\times K$ lead field matrix $\mathbf{L}$:
$\mathbf{L}\,\mathbf{y}=\mathbf{x},$ (1)
where $\mathbf{y}\in\mathbb{R}^{K}$ denotes the applied current pattern across $K$ electrode contacts, and $\mathbf{x}\in\mathbb{R}^{N+M}$ is the discretized volumetric current density with each entry corresponding to a local source element parameterized by three Cartesian dipolar degrees of freedom (DOFs), representing current density along the $x$-, $y$-, and $z$-axes. For our optimization framework, $\mathbf{L}$ is partitioned into target and nuisance components:
$\mathbf{L}=\begin{pmatrix}\mathbf{L}_{1}\\ \mathbf{L}_{2}\end{pmatrix},\qquad\mathbf{x}=\begin{pmatrix}\mathbf{x}_{1}\\ \mathbf{0}\end{pmatrix}$, (2)
where $\mathbf{L}_{1}\in\mathbb{R}^{N\times K}$ models the target region, $\mathbf{L}_{2}\in\mathbb{R}^{M\times K}$ models the nuisance region, and $\mathbf{x}_{1}\in\mathbb{R}^{N}$ defines the desired field in the target region. To ensure that only the direction of $\mathbf{x}_{1}$ is prescribed, we introduce the projection
$\mathbf{P}=\frac{\mathbf{x}_{1}\mathbf{x}_{1}^{\mathsf{T}}}{\|\mathbf{x}_{1}\|_{2}^{2}}\in\mathbb{R}^{N\times N},$ (3)
which enforces alignment of the induced field with $\mathbf{x}_{1}$. The resulting reduced formulation is:
$\begin{pmatrix}\mathbf{P}\mathbf{L}_{1}\\ \mathbf{L}_{2}\end{pmatrix}\mathbf{y}=\begin{pmatrix}\mathbf{P}\mathbf{x}_{1}\\ \mathbf{0}\end{pmatrix}$. (4)
The first block enforces field orientation within the target region, while the second enforces suppression outside it.
A.2 L1-norm Regularized L1-norm Fitting (L1L1)
The general form of the L1-norm regularized L1-norm fitting (L1L1) method described by Galaz Prieto et. al [galazprieto_2022_L1vsL2] is a linear programming optimization problem designed to minimize a linear objective function of continuous real variables subject to linear constraints. The optimization problem needs to find the best matching between $\bf y$}, and the focused field via ${\bf L}{\bf y}={\bf x}$. The task is
$\underset{y}{\textrm{minimize}}$ $\quad\left\|\begin{pmatrix}{\bf L}_{1}{\bf y}-{\bf x}_{1}\\ \Psi_{\varepsilon}[\nu^{-1}{\bf L}_{2}{\bf y}]\end{pmatrix}\right\|_{1}+{\alpha}\zeta\|{\bf y}\|_{1},$ (5)
subject to $\quad{\bf y}\preceq\gamma{\bf 1},$ (6)
$\quad\|{\bf y}\|_{1}\leq\mu,$ (7)
$\quad\sum_{\ell=1}^{K}y_{\ell}=0,\,\quad\ell=1,\ldots,K.$ (8)
The linear constraint 6 implies that the individual current injection for each $\ell$-th electrode is limited to a $\gamma$ value; constraint 7 implies the total current injection flowing through the system is within a safety $\mu$-value limit; and constraint 8 indicates the total sum of electrical activity from every active electrode in the system must be equal to zero. The regularization parameter $\bf\alpha$} sets the L1-regularization level regarding a scaling value $\zeta=\|{\bf L}\|_{1}$. The function
$\Psi_{\varepsilon}[{\bf w}]_{m}=\max\{\,|w_{m}|,\varepsilon\,\}\quad\hbox{for}\quad m=1,\ldots,M,$ (9)
where ${\bf w}=(w_{1},w_{2},\ldots,w_{M})$ sets the nuisance field threshold $0\leq\varepsilon\leq 1$ with respect to the scaling value $\nu=\|{\bf x}\|_{\infty}$, meaning that entries $({\bf L}_{2}{\bf y})_{m}$ with an absolute value below $\varepsilon\nu$ do not actively contribute to the minimization process due to the threshold. The constraint support index $\{\,m\,:\,|({\bf L}_{2}{\bf y})_{m}|\geq{\varepsilon\nu}\,\}$, i.e., sets the contributing to the value of the objective function.
To evaluate stimulation performance and to quantify the trade-off between target and undesired current spread, we define three decision variables: the focused current density $\Gamma$, the nuisance current density $\Xi$, and the resulting field ratio $\Theta$.
The focused current density $\Gamma$ (A/m2) quantifies the degree of field alignment between the induced current distribution and the desired stimulation target. It is defined as
$\Gamma=\frac{{\bf x}_{1}^{T}{\bf L}_{1}{\bf y}}{\|{\bf x}_{1}\|_{2}},$ (10)
where ${\bf x}_{1}\in\mathbb{R}^{N}$ represents the target volumetric current density template specifying the spatial and directional characteristics of the desired field within the stimulation region. The inner product ${\bf x}_{1}^{T}{\bf L}_{1}{\bf y}$ measures how strongly the induced field ${\bf L}_{1}{\bf y}$ aligns with the target pattern, while normalization by $\|{\bf x}_{1}\|_{2}$ scales the result to yield a field intensity expressed in A/m2.
The nuisance current density $\Xi$ (A/m2) characterizes the average magnitude of the induced field in off-target regions and is defined by
$\Xi=\frac{\|{\bf L}_{2}{\bf y}\|_{2}}{\sqrt{M}}.$ (11)
Dividing the Euclidean norm $\|{\bf L}_{2}{\bf y}\|_{2}$ by $\sqrt{M}$ yields the root-mean-square (RMS) field amplitude per node,
$\frac{\|{\bf L}_{2}{\bf y}\|_{2}}{\sqrt{M}}=\sqrt{\frac{1}{M}\sum_{m=1}^{M}\big(({\bf L}_{2}{\bf y})_{m}\big)^{2}}.$
The division by $\sqrt{M}$ converts the Euclidean norm into the root-mean-square (RMS) field amplitude per node, thereby making $\Xi$ invariant to the total number of discretization points in the nuisance domain. In cases where only a subset of nodes exhibits appreciable current density, the normalization may alternatively be restricted to the active constraint support, i.e., replacing $M$ with the number of nodes satisfying $|({\bf L}_{2}{\bf y})_{m}|\geq\varepsilon\nu$.
The field ratio $\Theta$ is introduced to characterize the overall selectivity of stimulation:
$\Theta=\frac{\Gamma}{\Xi}.$ (12)
A high $\Theta$ value reflects both strong current focusing within the target and minimal activation of non-target regions.
A.3 Tikhonov-Regularized Least Squares
TLS estimation [dmochowski2011optimized] is formulated as the following optimization problem:
$\min_{{\bf y}}\left\{\|{\bf L}_{1}{\bf y}-{\bf x}_{1}\|^{2}_{2}+\gamma^{2}\beta^{2}\|{\bf L}_{2}{\bf y}\|^{2}_{2}+\gamma^{2}\varsigma^{2}\|{\bf y}\|^{2}_{2}\right\},$ (13)
where $\varsigma=\|{\bf L}\|_{2}$. To enhance focality, the nuisance field weight $\beta\geq 0$ is treated as a variable parameter. The solution to (13) satisfies the linear system:
$\left({\bf L}_{1}^{T}{\bf L}_{1}+\gamma^{2}\beta^{2}{\bf L}_{2}^{T}{\bf L}_{2}+\gamma^{2}\varsigma^{2}{\bf I}\right){\bf y}={\bf L}_{1}^{T}{\bf x}_{1}.$ (14)
Both the regularization parameter and the nuisance field weight influence the above equations.